A copper wire is stretched by 1% office length calculate energy stored per unit volume of wire .(Y=[1.1.*10*-11 N/m2]
1
Expert's answer
2018-09-24T12:15:08-0400
Energy stored per unit volume is

E=1/2 s^2/Y=1/2 〖0.01〗^2/(1.1∙〖10〗^(-11) )=4.5∙〖10〗^6 J/m^3 .